HETAV PATEL

Grade 11,

mpare stability of O - 2 and O 2 + on basis of molecular orbit theory

mpare stability of O-2 and O2 on basis of molecular orbit theory

Grade:11

2 Answers

Arun
4 years ago
O2+ is more stable than O2 as bond order of o2+ is 2.5 and that of o2 is 2.
stability is directly propositional to bond order.
Vikas TU
4 years ago
Dear student 
Bond order of O2+ = (10-5 )/2 = 2.50 
Bond order of O2-= 10-7 /2 = 1.50 
Bond order of O2+ is greater than O2- 
So, O2+ will be more stable 
O2+ =  1s2 2s2 2p3 - Paramagnetic 
O2 - = 1s2 1s2 2p4 - Paramagnetic 
Hope this helps